o d curendar year unless otherwise statet 1. Your company uses the DDB method. Assets purchased betweeththey were depreciated for the entire month; assets purchased after the 15m as acquired the ollowing month. On May 1,20X1, it purchases for s253 with a useful life of 12 years and a salvage value of $10,000. book value at the end of 20X2? 1st and 15th are though they were he machines $250,000 a machi What is the machines ne a. $173,611 b. $185,185 c. $208,323 d. $215,.278 2 Your company uses the DDB method. Assets purchased between the 1t an depreciated for the entire month; assets purchased after the 15th as though they were acquired the following month. On July 1, 20X1, it purchases for $200,000 machine with a useful life of 10 years and a salvage value of $20,000. What is 20X2 depreciation expense? a. $38,800 b. $32,400 c. $32,000 d. $36,000 On January 1, 20X1, your company purchases for $500,000 a machine with a useful life of 12 years and a salvage value of $25,000. If your company uses DDB depreciation, what is 2004 expense? a. $48,225 b. $45,814 C. $47,839 d. $45,428 3. Your company uses the DDB method. Assets purchased between the 1st and 15th are depreciated for the entire month; assets purchased after the 15th as though they were acquired the following month. On April 6, 20X1, your firm purchases for $250,000 a machine that management estimates will last 15 years and have a salvage value of $10,000. What is 20X1 depreciation expense? a. $21,333 b. $25,000 c. $33,333 4. d. $32,000 Your firm uses 150% depreciation. Assets purchased between the 1st and 15th are depreciated for the entire month; assets purchased after the 15th as though they were acquired the following month. On March 12, 20X1, your company purchases for $200,000 a machine that it estimates will last 10 years and have a salvage value of 5. $60,000. What is 20X1 depreciation expense? a. $21,000 b. $30,000 C. $17,500 d. $25,000
Answer #1

Ans-1-Straight line rate= 1/ 12* 100= 8.33%

DDB rate= 2*8.3333= 16.6667%

20x1- $250,000*16.6667% *8/12 months= $27,777.8333

Book Value at the end of 20x1= $250,000-$27,777.8333= $222,222.1667

20x2- $222,222.1667* 16.6667%= $37,037.10

Book Value at the end of 20x2- $222,222.1667-$37,037.10= $185,185

Hence the correct option is b- $185,185

Ans-2 Straight line rate = 1/10* 100= 10%

DDB rate= 2 *10%= 20%

20x1- $200,000 *20% *6/12 months= 20,000

Book value at the end of 20x1= $200,000-$20,000= $180,000

20x2- $180,000 *20%= $36,000

20x2 depreciation expenses- $36,000

Hence, the correct option is d-$36,000

Ans-3-Straight line rate= 1/12*100= 8.333%

DDB rate= 2*8.333% = 16.6667%

Year Depreciation Expenses Accumulated Depreciation Book Value at the end of year
20x1

$83,333

($500,000*16.667%)

$83,333

$416,667

($500,000-$83,333)

20x2

$69,444

($416,667*16.6667%)

$152,777

($69,444+$83,333)

$347,223

($416,667-$69,444)

20x3

$57,870

($347,223* 16.6667%)

$210,647

($152,777+$57,870)

$289,353

($347,223- $57,870)

20x4

$48,225

($289,353*16.6667%)

Hence, the correct option is a. $48,225

Ans-4- Straight line rate= 1/15*100= 6.6667%

DDB rate= 2* 6.6667%= 13.3333%

A machine purchases $250,000 on april 6, 20x1

20x1 depreciation expenses:- $250,000* 13.3333% *9/12 months= $25,000

Hence, the correct option is b- $25,000

Ans-5- Useful life= 10 years, so the straight line depreciation rate= 1/10*100= 10%

DDB rate= 10%* 150%= 10%*1.5= 15%

Depreciation expense for 20x1 is= $200,000*15%* 10/12 months= $25,000

Hence, the correct option is d. $25,000